# Shuffle Products, Degenerate Affine Hecke Algebras, and Quantum Toda Lattice

### v2

@@ 7.1. Comparison with geometry
Observe that the composition of Theorem˜5.9 and Proposition˜7.13 gives an algebra map
$\mathrm{Y}_{-N,N}(\mathfrak{gl}_{2})^{\geq 0}\rightarrow\mathrm{U}_{\hbar}(\mathfrak{gl}^{R}_{N})\ltimes S^{\bullet}\mathbf{C}^{N}/\!\!/_{\psi}\,N_{+}^{R}\rightarrow N^{L}_{+}\,\backslash\!\!\backslash_{-\psi}\,\mathrm{D}_{\hbar}(\mathrm{GL}_{N})/\!\!/_{\psi}\,N_{+}^{R}\rightarrow H_{\bullet}^{\mathrm{GL}_{N}[[t]]\rtimes\mathbf{C}^{*}}(\mathrm{Gr}),$
@@
Likewise, the dual versions of Theorem˜B.3 and Proposition˜B.6 give
$\mathrm{Y}_{-N,N}(\mathfrak{gl}_{2})^{\leq 0}\rightarrow\mathrm{U}_{\hbar}(\mathfrak{gl}^{R}_{N})\ltimes S^{\bullet}(\mathbf{C}^{N})^{*}/\!\!/_{\psi}\,N_{+}^{R}\rightarrow N^{L}_{+}\,\backslash\!\!\backslash_{-\psi}\,\mathrm{D}_{\hbar}(\mathrm{GL}_{N})/\!\!/_{\psi}\,N_{+}^{R}\rightarrow H_{\bullet}^{\mathrm{GL}_{N}[[t]]\rtimes\mathbf{C}^{*}}(\mathrm{Gr}),$
$d_{1}(u)\mapsto A(u),\qquad x_{-}(u)\mapsto\frac{1}{u-c_{1}^{\mathrm{GL}_{N}}(\mathcal{Q})-\hbar}\cap[\mathrm{Gr}^{\omega_{1}^{*}}].$
By Theorem˜7.9, the combination of algebraic maps gives an isomorphism
@@
The explicit identification (7.4) of [BF08] comes from two facts. On the algebraic side, in loc. cit., the authors identify the center $\mathbf{Z}_{\mathfrak{gl}_{N}}$ with $W$-invariant functions on $\mathfrak{h}_{N}^{*}$ shifted by $\hbar\rho_{N}$ via the projection
@@
for instance, see [BF08, Lemma 5]. By definition of the quantum minor (1.9), we have
$T_{1\ldots N}^{1\ldots N}(u)\mapsto(u+E_{11}-N\hbar+\hbar)\ldots(u+E_{N-1,N-1}-\hbar)(u+E_{NN}),$
@@
7.2. Monopole operators
Recall the notations from Section˜1.1. Let $\lambda$ be a dominant weight, and denote by $v^{*}_{\lambda}\in V_{\lambda}^{*}$ the vector dual to the lowest-weight vector in the representation $V_{\lambda}$. Observe that the function $\langle g\cdot v_{\lambda},v^{*}_{\lambda}\rangle$ is $N_{+}^{L}\times N_{+}^{R}$-invariant, hence defines an element in the Toda lattice $N^{L}_{+}\,\backslash\!\!\backslash_{-\psi}\,\mathrm{D}_{\hbar}(\mathrm{GL}_{N})/\!\!/_{\psi}\,N_{+}^{R}$. The goal of this subsection is to compute these elements and to identify their images under the GKLO map for fundamental weights. This is an algebraic counterpart of the simplest case of the monopole operators formula [BFN19, (A.5)].
Under the embedding $V_{\omega_{k}}=\Lambda^{i}\mathbf{C}^{N}\hookrightarrow(\mathbf{C}^{N})^{\otimes k}$, let us choose a highest-weight vector and the dual of a lowest-weight vector as
$v_{\omega_{i}}=\sum_{\sigma\in S_{k}}(-1)^{\sigma}v_{\sigma(k)}\otimes\ldots\otimes v_{\sigma(1)},\qquad v^{*}_{\omega_{i}}=\phi_{N}\wedge\ldots\wedge\phi_{N-k+1}.$
@@ 7.2. Monopole operators
Under the embedding $^{R}\iota_{N}$ of Example˜3.5, we have
$\langle g\cdot v_{\omega_{k}},v^{*}_{\omega_{k}}\rangle={}^{R}\iota_{N}\left[\frac{1}{k!}\left(\prod_{1\leq i\neq j\leq k}(\Omega_{i}-\Omega_{j}-\hbar)\right)(v_{1})^{\otimes k}\right].$
@@
Observe that
$\langle g\cdot v_{\omega_{k}},v^{*}_{\omega_{k}}\rangle=k!\langle g\cdot v_{k}\otimes\ldots\otimes v_{1},\phi_{N}\wedge\ldots\wedge\phi_{N-k+1}\rangle.$
Consider the map
$^{L}\iota^{*}_{k}\otimes\ldots\otimes{}^{L}\iota^{*}_{1}\colon\mathrm{U}_{\hbar}(\mathfrak{gl}^{L}_{N})\otimes((\mathbf{C}^{N})^{*})^{\otimes k}\rightarrow\mathrm{D}_{\hbar}(\mathrm{GL}_{N}),$
@@
$(\mathrm{U}_{\hbar}(\mathfrak{gl}_{N}^{L})\otimes(\mathbf{C}^{N})^{*})\otimes_{\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})}\ldots\otimes_{\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})}(\mathrm{U}_{\hbar}(\mathfrak{gl}_{N}^{L})\otimes(\mathbf{C}^{N})^{*})\xrightarrow{{}^{L}\iota^{*}_{k}\otimes\ldots\otimes{}^{L}\iota^{*}_{1}}$
$\rightarrow\mathrm{D}_{\hbar}(\mathrm{GL}_{N})\otimes_{\mathrm{U}_{\hbar}(\mathfrak{gl}_{N}^{L})}\ldots\otimes_{\mathrm{U}_{\hbar}(\mathfrak{gl}_{N}^{L})}\mathrm{D}_{\hbar}(\mathrm{GL}_{N})\xrightarrow{m}\mathrm{D}_{\hbar}(\mathrm{GL}_{N})$
@@
$\mathrm{U}_{\hbar}(\mathfrak{gl}_{N}^{L})\otimes((\mathbf{C}^{N})^{*})^{\otimes k}\cong(\mathrm{U}_{\hbar}(\mathfrak{gl}_{N}^{L})\otimes(\mathbf{C}^{N})^{*})\otimes_{\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})}\ldots\otimes_{\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})}(\mathrm{U}_{\hbar}(\mathfrak{gl}_{N}^{L})\otimes(\mathbf{C}^{N})^{*}).$
In what follows, we denote by the same letter the induced map of the Kostant-Whittaker reductions
$^{L}\iota^{*}_{k}\otimes\ldots\otimes{}^{L}\iota^{*}_{1}\colon N^{L}_{+}\,\backslash\!\!\backslash_{-\psi}\,\mathrm{U}_{\hbar}(\mathfrak{gl}^{L}_{N})\otimes((\mathbf{C}^{N})^{*})^{\otimes k}\rightarrow N^{L}_{+}\,\backslash\!\!\backslash_{-\psi}\,\mathrm{D}_{\hbar}(\mathrm{GL}_{N}),$
@@
$\langle g\cdot v_{k}\otimes\ldots\otimes v_{1},\phi_{N-k+1}\wedge\ldots\wedge\phi_{N}\rangle=({}^{L}\iota^{*}_{k}\otimes\ldots\otimes{}^{L}\iota^{*}_{1})(\phi_{N}\wedge\ldots\wedge\phi_{N-k+1}),$
By Proposition˜B.5,
$({}^{L}\iota^{*}_{k}\otimes\ldots\otimes{}^{L}\iota^{*}_{1})(\phi_{N}\wedge\ldots\wedge\phi_{N-k+1})=$
$=$ $({}^{L}\iota^{*}_{k}\otimes\ldots\otimes{}^{L}\iota^{*}_{1})\left[\frac{1}{k!}\prod_{i<j}(\Omega^{*}_{i}-\Omega^{*}_{j}+\hbar)\cdot(\phi_{N}^{\otimes k})\right]\in N^{L}_{+}\,\backslash\!\!\backslash_{-\psi}\,\mathrm{D}_{\hbar}(\mathrm{GL}_{N})$
(recall that $\psi=1$). At the same time, it follows from Proposition˜7.2 and (1.14) that
$^{L}\iota_{\alpha}^{*}[\Omega^{*}(\phi_{N})]=[XE^{L}]_{\alpha N}=[(-E^{R}+N\hbar)X]_{\alpha N}={}^{R}\iota_{N}[(-\Omega+N\hbar)(v_{\alpha})]$
@@
$({}^{L}\iota^{*}_{k}\otimes\ldots\otimes{}^{L}\iota^{*}_{1})\left[\frac{1}{k!}\prod_{i<j}(\Omega^{*}_{i}-\Omega^{*}_{j}+\hbar)\cdot(\phi_{N}^{\otimes k})\right]=$
$=$ $(-1)^{\frac{k(k-1)}{2}}({}^{R}\iota_{N}\otimes\ldots\otimes{}^{R}\iota_{N})\left[\frac{1}{k!}\prod_{i<j}(\Omega_{i}-\Omega_{j}-\hbar)\cdot(v_{k}\otimes\ldots\otimes v_{1})\right],$
@@
$^{R}\iota_{N}\otimes\ldots\otimes{}^{R}\iota_{N}\colon\mathrm{U}_{\hbar}(\mathfrak{gl}^{R}_{N})\otimes(\mathbf{C}^{N})^{\otimes k}\rightarrow N^{L}_{+}\,\backslash\!\!\backslash_{-\psi}\,\mathrm{D}_{\hbar}(\mathrm{GL}_{N})$
@@
$({}^{R}\iota_{N}\otimes\ldots\otimes{}^{R}\iota_{N})\left[\prod_{i<j}(\Omega_{i}-\Omega_{j}-\hbar)\cdot(v_{k}\otimes\ldots\otimes v_{1})\right]=\langle g\cdot v_{\omega_{k}},v^{*}_{\omega_{k}}\rangle\in N^{L}_{+}\,\backslash\!\!\backslash_{-\psi}\,\mathrm{D}_{\hbar}(\mathrm{GL}_{N})/\!\!/_{\psi}\,N_{+}^{R},$
@@
$({}^{R}\iota_{N}\otimes\ldots\otimes{}^{R}\iota_{N})\left[\prod_{i<j}(\Omega_{i}-\Omega_{j}-\hbar)\cdot(v_{k}\otimes\ldots\otimes v_{1})\right]=({}^{R}\iota_{N}\otimes\ldots\otimes{}^{R}\iota_{N})\left[\prod_{i<j}(\Omega_{i}-\Omega_{j}-\hbar)\cdot P_{\psi}(v_{k}\otimes\ldots\otimes v_{1})P_{\psi}\right],$
where the argument of the right-hand side is considered as an element of $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N}^{R})\otimes(\mathbf{C}^{N})^{\otimes k}/\!\!/_{\psi}\,N_{+}^{R}$. We have a commutative triangle
${\mathrm{U}_{\hbar}(\mathfrak{gl}^{R}_{N})\otimes(\mathbf{C}^{N})^{\otimes k}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{{}^{R}\iota_{N}\otimes\ldots\otimes{}^{R}\iota_{N}}$$\scriptstyle{e_{k}}$${N^{L}_{+}\,\backslash\!\!\backslash_{-\psi}\,\mathrm{D}_{\hbar}(\mathrm{GL}_{N})}$${\mathrm{U}_{\hbar}(\mathfrak{gl}^{R}_{N})\otimes S^{\bullet}\mathbf{C}^{N}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{{}^{R}\iota_{N}}$
where $e_{k}\colon(\mathbf{C}^{N})^{\otimes k}\rightarrow S^{k}\mathbf{C}^{N}$ is the symmetrizer map. Also, by Proposition˜5.10, we have
$P_{\psi}^{R}v_{k}\otimes\ldots\otimes v_{1}P_{\psi}^{R}=\Omega_{1}^{k-1}\ldots\Omega_{k-1}^{1}\Omega_{k}^{0}(v_{1}^{\otimes k})P_{\psi}^{R}\in\mathrm{U}_{\hbar}(\mathfrak{gl}_{N}^{R})\otimes(\mathbf{C}^{N})^{\otimes k}/\!\!/_{\psi}\,N_{+}^{R}.$
@@
$(-1)^{\frac{k(k-1)}{2}}{}^{R}\iota_{N}\left[e_{k}\cdot\prod_{i<j}(\Omega_{i}-\Omega_{j}-\hbar)\Omega_{1}^{k-1}\ldots\Omega_{k-1}^{1}\Omega_{k}^{(}v_{1}^{\otimes k})\right]\in\mathrm{U}_{\hbar}(\mathfrak{gl}_{N}^{R})\otimes(\mathbf{C}^{N})^{\otimes k}/\!\!/_{\psi}\,N^{R}_{+},$
which, by Proposition˜5.7, is given by the action of the element
$(-1)^{\frac{k(k-1)}{2}}\left(e^{\hbar}_{k}\cdot\prod_{i<j}(Y_{i}-Y_{j}-\hbar)Y_{1}^{k-1}\ldots Y_{k-1}^{1}Y_{k}^{0}\cdot e^{\hbar}_{k}\right)\in\mathbf{C}[\hbar][Y_{1},\ldots,Y_{k}]^{S_{k}}$
in the spherical subalgebra of $\mathfrak{H}^{\hbar}_{k}$ as in (2.7). By Proposition˜2.7, it is equal to
$\frac{(-1)^{\frac{k(k-1)}{2}}}{k!}\mathrm{Sym}_{k}\left[\prod_{i<j}\frac{(Y_{i}-Y_{j}+\hbar)(Y_{i}-Y_{j}-\hbar)}{Y_{i}-Y_{j}}Y_{1}^{k-1}\ldots Y_{k-1}\right]=$
$=$ $\frac{(-1)^{\frac{k(k-1)}{2}}}{k!}\prod_{i<j}(Y_{i}-Y_{j}+\hbar)(Y_{i}-Y_{j}-\hbar)\cdot\mathrm{Sym}_{k}\left[\prod_{i<j}(Y_{i}-Y_{j})Y_{1}^{k-1}\ldots Y_{k-1}^{1}Y_{k}^{0}\right]$
@@
$\mathrm{Sym}_{k}\left[\prod_{i<j}(Y_{i}-Y_{j})Y_{1}^{k-1}\ldots Y_{k-1}^{1}Y_{k}^{0}\right]=\prod_{i<j}(Y_{i}-Y_{j})\sum_{\sigma\in S_{k}}(-1)^{\sigma}Y_{\sigma(1)}^{k-1}\ldots Y_{\sigma(k-1)}^{1}Y_{\sigma(k)}^{0}=1.$
@@
$\langle g\cdot v_{\omega_{i}},v^{*}_{\omega_{i}}\rangle={}^{R}\iota_{N}\left[\frac{1}{k!}\left(\prod_{1\leq i\neq j\leq k}(\Omega_{i}-\Omega_{j}-\hbar)\right)(v_{1})^{\otimes k}\right],$
@@
The next theorem computes the images of these elements under the GKLO map of Theorem˜7.9.
@@
of Theorem˜7.9, we have
$\langle g\cdot v_{\omega_{k}},v^{*}_{\omega_{k}}\rangle\mapsto\sum_{\begin{subarray}{c}I\subset\{1,\ldots,N\}\\ |I|=k\end{subarray}}\prod_{\begin{subarray}{c}i\in I\\ j\in\bar{I}\end{subarray}}\frac{1}{w_{i}-w_{j}}\prod_{i\in I}\mathbf{u}_{i}^{-1},$
where $\bar{I}$ is the complement of $I$ in $\{1,\ldots,N\}$.
@@
Since $\langle g\cdot v_{\omega_{k}},v^{*}_{\omega_{k}}\rangle$ is in the image of $^{R}\iota_{N}$, it is enough to prove the statement for the Miura transform
@@
of Theorem˜6.7. Recall that

### v4

@@ 8.1 Comparison with geometry
Observe that the composition of Theorem 6.9 and Proposition 8.14 gives an algebra map
$\mathrm{Y}_{-N,N}(\mathfrak{gl}_{2})^{\geq 0}\rightarrow\mathrm{U}_{\hbar}\bigl(\mathfrak{gl}^{R}_{N}\bigr)\ltimes S^{\bullet}\mathbf{C}^{N}/\!\!/_{\psi}\,N_{+}^{R}$
[…]
@@
Likewise, the dual versions of Theorem B.3 and Proposition B.6 give
$\mathrm{Y}_{-N,N}(\mathfrak{gl}_{2})^{\leq 0}\rightarrow\mathrm{U}_{\hbar}\bigl(\mathfrak{gl}^{R}_{N}\bigr)\ltimes S^{\bullet}\bigl(\mathbf{C}^{N}\bigr)^{*}/\!\!/_{\psi}\,N_{+}^{R}$
[…]
$d_{1}(u)\mapsto A(u),\qquad x_{-}(u)\mapsto\frac{1}{u-c_{1}^{\mathrm{GL}_{N}}(\mathcal{Q})-\hbar}\cap\bigl[\mathrm{Gr}^{\omega_{1}^{*}}\bigr].$
By Theorem 8.10, the combination of algebraic maps gives an isomorphism
@@
The explicit identification (8.5) of [3] comes from two facts. On the algebraic side, in loc. cit., the authors identify the center $\mathbf{Z}_{\mathfrak{gl}_{N}}$ with $W$-invariant functions on $\mathfrak{h}_{N}^{*}$ shifted by $\hbar\rho_{N}$ via the projection
@@
for instance, see [3, Lemma 5]. By definition of the quantum minor (2.7), we have
$T_{1\dots N}^{1\dots N}(u)\mapsto(u+E_{11}-N\hbar+\hbar)\cdots(u+E_{N-1,N-1}-\hbar)(u+E_{NN}),$
@@
8.2 Monopole operators
Recall the notations from Section 2.1. Let $\lambda$ be a dominant weight, and denote by $v^{*}_{\lambda}\in V_{\lambda}^{*}$} the vector dual to the lowest-weight vector in the representation $V_{\lambda}$. Observe that the function $\langle g\cdot v_{\lambda},v^{*}_{\lambda}\rangle$} is $N_{+}^{L}\times N_{+}^{R}$-invariant, and hence defines an element in the Toda lattice $N^{L}_{+}\,\backslash\!\!\backslash_{-\psi}\,\mathrm{D}_{\hbar}(\mathrm{GL}_{N})/\!\!/_{\psi}\,N_{+}^{R}$. The goal of this subsection is to compute these elements and to identify their images under the GKLO map for fundamental weights. This is an algebraic counterpart of the simplest case of the monopole operators formula [4, formula (A.5)].
Under the embedding $V_{\omega_{k}}=\Lambda^{i}\mathbf{C}^{N}\hookrightarrow\bigl(\mathbf{C}^{N}\bigr)^{\otimes k}$, let us choose a highest-weight vector and the dual of a lowest-weight vector as
$v_{\omega_{i}}=\sum_{\sigma\in S_{k}}(-1)^{\sigma}v_{\sigma(k)}\otimes\dots\otimes v_{\sigma(1)},\qquad v^{*}_{\omega_{i}}=\phi_{N}\wedge\dots\wedge\phi_{N-k+1}.$
@@ 8.2 Monopole operators
Under the embedding $^{R}\iota_{N}$ of Example 4.5, we have
$\bigl\langle g\cdot v_{\omega_{k}},v^{*}_{\omega_{k}}\bigr\rangle={}^{R}\iota_{N}\biggl[\frac{1}{k!}\biggl(\prod_{1\leq i\neq j\leq k}(\Omega_{i}-\Omega_{j}-\hbar)\biggr)(v_{1})^{\otimes k}\biggr].$
@@
Observe that $\bigl\langle g\cdot v_{\omega_{k}},v^{*}_{\omega_{k}}\bigr\rangle=k!\langle g\cdot v_{k}\otimes\dots\otimes v_{1},\phi_{N}\wedge\dots\wedge\phi_{N-k+1}\rangle$. Consider the map
$^{L}\iota^{*}_{k}\otimes\dots\otimes{}^{L}\iota^{*}_{1}\colon\ \mathrm{U}_{\hbar}\bigl(\mathfrak{gl}^{L}_{N}\bigr)\otimes\bigl(\bigl(\mathbf{C}^{N}\bigr)^{*}\bigr)^{\otimes k}\rightarrow\mathrm{D}_{\hbar}(\mathrm{GL}_{N}),$
@@
$\bigl(\mathrm{U}_{\hbar}\bigl(\mathfrak{gl}_{N}^{L}\bigr)\otimes\bigl(\mathbf{C}^{N}\bigr)^{*}\bigr)\otimes_{\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})}\dots\otimes_{\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})}\bigl(\mathrm{U}_{\hbar}\bigl(\mathfrak{gl}_{N}^{L}\bigr)\otimes\bigl(\mathbf{C}^{N}\bigr)^{*}\bigr)$
$\qquad\xrightarrow{{}^{L}\iota^{*}_{k}\otimes\dots\otimes{}^{L}\iota^{*}_{1}}\mathrm{D}_{\hbar}(\mathrm{GL}_{N})\otimes_{\mathrm{U}_{\hbar}(\mathfrak{gl}_{N}^{L})}\dots\otimes_{\mathrm{U}_{\hbar}(\mathfrak{gl}_{N}^{L})}\mathrm{D}_{\hbar}(\mathrm{GL}_{N})\xrightarrow{m}\mathrm{D}_{\hbar}(\mathrm{GL}_{N})$
@@
$\mathrm{U}_{\hbar}\bigl(\mathfrak{gl}_{N}^{L}\bigr)\otimes\bigl(\bigl(\mathbf{C}^{N}\bigr)^{*}\bigr)^{\otimes k}\cong\bigl(\mathrm{U}_{\hbar}\bigl(\mathfrak{gl}_{N}^{L}\bigr)\otimes\bigl(\mathbf{C}^{N}\bigr)^{*}\bigr)\otimes_{\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})}\dots\otimes_{\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})}\bigl(\mathrm{U}_{\hbar}\bigl(\mathfrak{gl}_{N}^{L}\bigr)\otimes\bigl(\mathbf{C}^{N}\bigr)^{*}\bigr).$
In what follows, we denote by the same letter the induced map of the Kostant–Whittaker reductions
$^{L}\iota^{*}_{k}\otimes\dots\otimes{}^{L}\iota^{*}_{1}\colon\ N^{L}_{+}\,\backslash\!\!\backslash_{-\psi}\,\mathrm{U}_{\hbar}\bigl(\mathfrak{gl}^{L}_{N}\bigr)\otimes\bigl(\bigl(\mathbf{C}^{N}\bigr)^{*}\bigr)^{\otimes k}\rightarrow N^{L}_{+}\,\backslash\!\!\backslash_{-\psi}\,\mathrm{D}_{\hbar}(\mathrm{GL}_{N}),$
@@
$\langle g\cdot v_{k}\otimes\dots\otimes v_{1},\phi_{N-k+1}\wedge\dots\wedge\phi_{N}\rangle=\bigl({}^{L}\iota^{*}_{k}\otimes\dots\otimes{}^{L}\iota^{*}_{1}\bigr)(\phi_{N}\wedge\dots\wedge\phi_{N-k+1}),$
By Proposition B.5,
$\bigl({}^{L}\iota^{*}_{k}\otimes\dots\otimes{}^{L}\iota^{*}_{1}\bigr)(\phi_{N}\wedge\dots\wedge\phi_{N-k+1})$
$\qquad{}=\bigl({}^{L}\iota^{*}_{k}\otimes\dots\otimes{}^{L}\iota^{*}_{1}\bigr)\biggl[\frac{1}{k!}\prod_{i<j}\bigl(\Omega^{*}_{i}-\Omega^{*}_{j}+\hbar\bigr)\cdot\bigl(\phi_{N}^{\otimes k}\bigr)\biggr]\in N^{L}_{+}\,\backslash\!\!\backslash_{-\psi}\,\mathrm{D}_{\hbar}(\mathrm{GL}_{N})$
(recall that $\psi=1$). At the same time, it follows from Proposition 8.3 and (2.11) that
$^{L}\iota_{\alpha}^{*}[\Omega^{*}(\phi_{N})]=\bigl[XE^{L}\bigr]_{\alpha N}=\bigl[\bigl(-E^{R}+N\hbar\bigr)X\bigr]_{\alpha N}={}^{R}\iota_{N}[(-\Omega+N\hbar)(v_{\alpha})]$
@@
$\bigl({}^{L}\iota^{*}_{k}\otimes\dots\otimes{}^{L}\iota^{*}_{1}\bigr)\biggl[\frac{1}{k!}\prod_{i<j}\bigl(\Omega^{*}_{i}-\Omega^{*}_{j}+\hbar\bigr)\cdot\bigl(\phi_{N}^{\otimes k}\bigr)\biggr]$
$\qquad{}=(-1)^{\frac{k(k-1)}{2}}\bigl({}^{R}\iota_{N}\otimes\dots\otimes{}^{R}\iota_{N}\bigr)\biggl[\frac{1}{k!}\prod_{i<j}(\Omega_{i}-\Omega_{j}-\hbar)\cdot(v_{k}\otimes\dots\otimes v_{1})\biggr],$
@@
$^{R}\iota_{N}\otimes\dots\otimes{}^{R}\iota_{N}\colon\ \mathrm{U}_{\hbar}\bigl(\mathfrak{gl}^{R}_{N}\bigr)\otimes\bigl(\mathbf{C}^{N}\bigr)^{\otimes k}\rightarrow N^{L}_{+}\,\backslash\!\!\backslash_{-\psi}\,\mathrm{D}_{\hbar}(\mathrm{GL}_{N})$
@@
$\bigl({}^{R}\iota_{N}\otimes\dots\otimes{}^{R}\iota_{N}\bigr)\biggl[\prod_{i<j}(\Omega_{i}-\Omega_{j}-\hbar)\cdot(v_{k}\otimes\dots\otimes v_{1})\biggr]$
$\qquad{}=\bigl\langle g\cdot v_{\omega_{k}},v^{*}_{\omega_{k}}\bigr\rangle\in N^{L}_{+}\,\backslash\!\!\backslash_{-\psi}\,\mathrm{D}_{\hbar}(\mathrm{GL}_{N})/\!\!/_{\psi}\,N_{+}^{R},$
@@
$\bigl({}^{R}\iota_{N}\otimes\dots\otimes{}^{R}\iota_{N}\bigr)\biggl[\prod_{i<j}(\Omega_{i}-\Omega_{j}-\hbar)\cdot(v_{k}\otimes\dots\otimes v_{1})\biggr]$
$\qquad{}=\bigl({}^{R}\iota_{N}\otimes\dots\otimes{}^{R}\iota_{N}\bigr)\biggl[\prod_{i<j}(\Omega_{i}-\Omega_{j}-\hbar)\cdot P_{\psi}(v_{k}\otimes\dots\otimes v_{1})P_{\psi}\biggr]$
(the argument of the right-hand side is considered as an element of $\mathrm{U}_{\hbar}\bigl(\mathfrak{gl}_{N}^{R}\bigr)\otimes\bigl(\mathbf{C}^{N}\bigr)^{\otimes k}/\!\!/_{\psi}\,N_{+}^{R}$). We have a commutative triangle
${\mathrm{U}_{\hbar}\bigl(\mathfrak{gl}^{R}_{N}\bigr)\otimes\bigl(\mathbf{C}^{N}\bigr)^{\otimes k}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{{}^{R}\iota_{N}\otimes\dots\otimes{}^{R}\iota_{N}}$$\scriptstyle{e_{k}}$${N^{L}_{+}\,\backslash\!\!\backslash_{-\psi}\,\mathrm{D}_{\hbar}(\mathrm{GL}_{N})}$${\mathrm{U}_{\hbar}\bigl(\mathfrak{gl}^{R}_{N}\bigr)\otimes S^{\bullet}\mathbf{C}^{N}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{{}^{R}\iota_{N}}$
where $e_{k}\colon\bigl(\mathbf{C}^{N}\bigr)^{\otimes k}\rightarrow S^{k}\mathbf{C}^{N}$ is the symmetrizer map. Also, by Proposition 6.10, we have
$P_{\psi}^{R}v_{k}\otimes\dots\otimes v_{1}P_{\psi}^{R}=\Omega_{1}^{k-1}\cdots\Omega_{k-1}^{1}\Omega_{k}^{0}\bigl(v_{1}^{\otimes k}\bigr)P_{\psi}^{R}\in\mathrm{U}_{\hbar}\bigl(\mathfrak{gl}_{N}^{R}\bigr)\otimes\bigl(\mathbf{C}^{N}\bigr)^{\otimes k}/\!\!/_{\psi}\,N_{+}^{R}.$
@@
$(-1)^{\frac{k(k-1)}{2}}{}^{R}\iota_{N}\biggl[e_{k}\cdot\prod_{i<j}(\Omega_{i}-\Omega_{j}-\hbar)\Omega_{1}^{k-1}\cdots\Omega_{k-1}^{1}\Omega_{k}^{0}\bigl(v_{1}^{\otimes k}\bigr)\biggr]\in\mathrm{U}_{\hbar}\bigl(\mathfrak{gl}_{N}^{R}\bigr)\otimes\bigl(\mathbf{C}^{N}\bigr)^{\otimes k}/\!\!/_{\psi}\,N^{R}_{+},$
which, by Proposition 6.7, is given by the action of the element
$(-1)^{\frac{k(k-1)}{2}}\biggl({\rm e}^{\hbar}_{k}\cdot\prod_{i<j}(Y_{i}-Y_{j}-\hbar)Y_{1}^{k-1}\cdots Y_{k-1}^{1}Y_{k}^{0}\cdot{\rm e}^{\hbar}_{k}\biggr)\in\mathbf{C}[\hbar][Y_{1},\dots,Y_{k}]^{S_{k}}$
in the spherical subalgebra of $\mathfrak{H}^{\hbar}_{k}$ as in (3.7). By Proposition 3.7, it is equal to
$\frac{(-1)^{\frac{k(k-1)}{2}}}{k!}\mathrm{Sym}_{k}\biggl[\prod_{i<j}\frac{(Y_{i}-Y_{j}+\hbar)(Y_{i}-Y_{j}-\hbar)}{Y_{i}-Y_{j}}Y_{1}^{k-1}\cdots Y_{k-1}^{1}Y_{k}^{0}\biggr]$
$\qquad{}=\frac{(-1)^{\frac{k(k-1)}{2}}}{k!}\prod_{i<j}(Y_{i}-Y_{j}+\hbar)(Y_{i}-Y_{j}-\hbar)\cdot\mathrm{Sym}_{k}\biggl[\prod_{i<j}(Y_{i}-Y_{j})Y_{1}^{k-1}\cdots Y_{k-1}^{1}Y_{k}^{0}\biggr].$
@@
$\mathrm{Sym}_{k}\biggl[\prod_{i<j}(Y_{i}-Y_{j})Y_{1}^{k-1}\cdots Y_{k-1}^{1}Y_{k}^{0}\biggr]=\prod_{i<j}(Y_{i}-Y_{j})\sum_{\sigma\in S_{k}}(-1)^{\sigma}Y_{\sigma(1)}^{k-1}\cdots Y_{\sigma(k-1)}^{1}Y_{\sigma(k)}^{0}=1.$
@@
$\bigl\langle g\cdot v_{\omega_{i}},v^{*}_{\omega_{i}}\bigr\rangle={}^{R}\iota_{N}\biggl[\frac{1}{k!}\biggl(\prod_{1\leq i\neq j\leq k}(\Omega_{i}-\Omega_{j}-\hbar)\biggr)(v_{1})^{\otimes k}\biggr],$
@@
The next theorem computes the images of these elements under the GKLO map of Theorem 8.10.
@@
of Theorem 8.10, we have
$\bigl\langle g\cdot v_{\omega_{k}},v^{*}_{\omega_{k}}\bigr\rangle\mapsto\sum_{\begin{subarray}{c}I\subset\{1,\dots,N\},\\ |I|=k\end{subarray}}\prod_{\begin{subarray}{c}i\in I,\\ j\in\bar{I}\end{subarray}}\frac{1}{w_{i}-w_{j}}\prod_{i\in I}\mathbf{u}_{i}^{-1},$
where $\bar{I}$ is the complement of $I$ in $\{1,\dots,N\}$.
@@
Since $\bigl\langle g\cdot v_{\omega_{k}},v^{*}_{\omega_{k}}\bigr\rangle$ is in the image of $^{R}\iota_{N}$, it is enough to prove the statement for the Miura transform
@@
of Theorem 7.7. Recall that